OpenStudy (anonymous):

range:

OpenStudy (anonymous):

\[y=\frac{ 8 }{ (4+x ^{2}) }\] the book says tha range is \[0<y \le2\] how come?

OpenStudy (anonymous):

i tried to solve for x to find it but i got nothing..

OpenStudy (lgbasallote):

cross multiply.. \[\implies y(4 + x^2) = 8\] distribute \[\implies 4y + x^2y = 8\] isolate x \[\implies x^2y = 8 - 4y\] divide by y \[\implies x^2 = \frac{8-4y}{y}\] take the square root \[\implies x = \sqrt{\frac{8-4y}{y}}\] y cant be 0 because it will be undefined y cant be greater than 2 because the radicand will be negative so \[0<y\le 2\] does that help?

OpenStudy (anonymous):

yes, thank you! i see where i went wrong, i had changed the y to zero before multiplying :)

OpenStudy (lgbasallote):

i see. glad you saw your mistake